# ContrastScore: Towards Higher Quality, Less Biased, More Efficient Evaluation Metrics with Contrastive Evaluation

### v1

@@ 5.3 Efficiency Analysis
Llama Summ WMT Qwen Summ WMT
Single 1B 63.29 339.04 0.5B 141.96 552.68
3B 27.08 153.25 3B 27.69 194.22
8B 26.43 72.17 7B 26.43 99.84
Ensemble (3B,1B) 39.51 110.63 (3B,0.5B) 44.94 154.5
(8B,1B) 23.81 66.8 (7B,0.5B) 26.11 93.92
(8B,3B) 19.81 53.95 (7B,3B) 20.35 71.04
Contrast (3B,1B) 39.12 109.31 (3B,0.5B) 44.06 153.99
(8B,1B) 23.33 66.11 (7B,0.5B) 25.71 93.41
(8B,3B) 19.05 53.45 (7B,3B) 19.61 70.97
Table 5: Processing Speed for Machine Translation and summarization evaluation tasks. Measured in Samples Per Second with batch size of 16 on H100 GPUs in bfloat16 precision.
Figure 1: Performance and efficiency of ContrastScore with smaller models compared to single larger model on summarization task.
Table 5 reports processing speeds for both machine translation (WMT) and summarization (Summ) evaluation tasks, measured in samples per second with a batch size 16, running in bfloat16 precision. In the LLaMa family, the single large model LLaMa 8B processes 26.43 samples/s on Summ and 72.17 samples/s on WMT, whereas ContrastScore using two smaller models (LLaMa 3B, 1B) achieves 39.12 samples/s on Summ and 109.31 samples/s on WMT — about 1.5$\times$ faster in both cases. Similarly, in the Qwen family, the single large model Qwen 7B shows 26.43 samples/s on Summ and 99.84 samples/s on WMT, while ContrastScore with Qwen (3B, 0.5B) processes 44.06 samples/s on Summ and 153.99 samples/s on WMT (approximately 1.7$\times$ faster on Summ and 1.5$\times$ faster on WMT). Moreover, as shown in Tables 1(b) and 2, the ContrastScore method with two smaller models (for both model families) also produces higher correlation with human judgments compared to using a single larger model. Figure 1 shows that employing two smaller models in a ContrastScore framework yields both better efficiency and improved evaluation quality.
@@ 6 Further Analysis
Original CE-Score COH CON FL RE AVG
(3B,1B) 0.405 0.441 0.432 0.3373 0.404
Llama (8B,1B) 0.388 0.471 0.446 0.329 0.408
(8B,3B) 0.418 0.475 0.476 0.347 0.429
(3B,0.5B) 0.448 0.474 0.463 0.340 0.431
Qwen (7B,0.5B) 0.429 0.487 0.473 0.332 0.430
(7B,3B) 0.431 0.485 0.487 0.336 0.435
Table 6: Pearson correlation between scores based on the original ratio-based Contrastive Decoding formulation (Li et al., 2023) and human judgments, where the abbreviations COH, CON, FL, and RE stand for coherence, consistency, fluency, and relevance, respectively.
Source Text {CJK*}UTF8gbsn产品防伪标识在哪里
@@
Hypothesis 1 Human Rank: 1 Expert 0.2471 0.7305 0.9922 0.02881 0.9805 0.7969 1.000 0.000457 -0.717 2
@@
Hypothesis 2 Human Rank:2 Expert 0.2471 0.7305 0.9922 0.02881 0.9805 0.7969 1.000 0.002808 -0.618 1
@@
Hypothesis 3 Human Rank:3 Expert 0.005951 0.000168 0.2910 4.268e-05 0.001602 0.004944 -2.668 3
@@
Table 7: Case Study: Comparison of ContrastScore with a single expert model (Qwen 3B) and an amateur model (Qwen 0.5B) for Chinese-to-English (ZH-EN) machine translation. $\bar{\textbf{Log(P)}}$ denotes mean log-probability across all target tokens.
Weighted Ensemble.   We investigate whether the benefits of ContrastScore could be achieved through a simpler method based on probability combination. Specifically, we test a weighted ensemble approach that linearly combines the probabilities from both larger and smaller models.
@@
where $\gamma\in[0,1]$ is the weight factor that controls the contribution of each model. When $\gamma=0.5$, this formulation simplifies to the standard ensemble method, which averages probabilities as reported in our main results. To further analyze the impact of $\gamma$, we conduct a parameter sweep on the summarization task using Llama3.2 (3B, 1B), varying $\gamma$ from 0 to 1 with a step size of 0.05. The results, presented in Figure 3 in Appendix C, show that the weighted ensemble achieves its highest performance at $\gamma=0.55$, with an average correlation of 0.4694. However, $\gamma=0.5$ yields nearly identical results, suggesting that a simple averaging method is nearly as effective as the best-weighted combination. More importantly, ContrastScore (0.474) outperforms the optimal setting of weighted ensembling (0.469) in correlation, highlighting its advantage over probability ensembling approaches.
Ratio-based Contrastive Decoding Formulation. To demonstrate that ContrastScore’s discrepancy-based probability formulation is better suited for evaluation tasks than ratio-based contrast decoding formulation in Li et al. (2023), we test Eq. (4) on the summarization evaluation task.
Table 6 presents the Pearson correlation results between the ratio-based formulation in Li et al. (2023) and human scores for both the Llama and Qwen model families. In all cases, the ratio-based scores are inferior to single-model baselines. This can be attributed to the division formulation, where the score approaches 1 when both the expert and amateur models assign either high or low probabilities, thereby reducing the distinguishability for evaluation.
Case Study
To provide a qualitative analysis, Table 7 presents a case study comparing ContrastScore with a single expert model (Qwen 3B) and an amateur model (Qwen 0.5B) for Chinese-to-English (ZH-EN) machine translation. It can be observed that the expert model misranked Hypothesis 1 as the second-best translation, whereas human evaluators ranked it as the best (Rank 1). This misranking was largely influenced by the expert model assigning a very low probability (0.000457) to “logo”, a key token in the translation. In contrast, the amateur model assigned a significantly higher probability (0.06592) to “logo”, demonstrating greater confidence in its relevance. ContrastScore leveraged this probability discrepancy to adjust the probability of “logo”, thereby increasing the overall score of Hypothesis 1 and correctly ranking it as the best translation. This case illustrates how ContrastScore mitigates expert model underestimation of critical tokens, leading to improved evaluation alignment with human judgment.
In the case of Hypothesis 3, all methods (i.e., expert, amateur, and ContrastScore) correctly ranked it as the worst translation (Rank 3). This consistency occurred because both the expert and amateur models assigned low probabilities across all tokens, indicating poor translation quality. Unlike in Hypothesis 1, where significant probability discrepancies required adjustment, ContrastScore made minimal modifications since the expert and amateur models were already in agreement in their evaluation. This demonstrates that ContrastScore does not introduce unnecessary changes when the models agree, ensuring that it only refines scores when meaningful corrections are needed.
@@ 7 Conclusion
In this paper, we introduce ContrastScore, an evaluation metric that leverages structured disagreement between two language models of differing capacities. Unlike conventional LLM-based evaluation methods that rely solely on a single model’s likelihood estimates, ContrastScore incorporates a weaker auxiliary model as a contrastive signal to adjust probability scores for better alignment with human judgments. Extensive experiments demonstrate that ContrastScore consistently achieves stronger correlation with human evaluations than both single-model and ensemble-based baselines. Furthermore, ContrastScore is efficient and can effectively mitigate common evaluation biases, such as preferences for length and likelihood, resulting in more robust evaluation.
Appendix A Prompts
The prompts for the summarization and machine translation tasks in source-based setting are presented in Table 8. For the reference-based setting, the prompts are "To rephrase it" and "That is to say".
@@ Appendix A Prompts
Appendix B The impact of hyperparameter gamma
@@
Figure 2: Exploration of the impacts of $\gamma$. Testing correlation between evaluator score and human score in ZH-EN language pair of machine translation task.
@@ Appendix C Results
Figure 4: Correlation of various evaluation scores with human score on WMT dataset
Figure 5: Correlation of various evaluation scores with human score on SummEval dataset

### v2

@@ 5.3 Efficiency Analysis
Figure 1: Efficiency and quality of ContrastScore with smaller models compared to single larger model on summarization task.
[…]
@@ 6 Further Analysis
Source Text 产品防伪标识在哪里
@@
Hypothesis 1
Human Rank: 1
Expert 0.2471 0.7305 0.9922 0.02881 0.9805 0.7969 1.000 0.000457 -0.717 2
@@
Hypothesis 2
Human Rank:2
Expert 0.2471 0.7305 0.9922 0.02881 0.9805 0.7969 1.000 0.002808 -0.618 1
@@
Hypothesis 3
Human Rank:3
Expert 0.005951 0.000168 0.2910 4.268e-05 0.001602 0.004944 -2.668 3
@@
Table 5: Case Study: Comparison of ContrastScore with Qwen 3B as an expert and Qwen 0.5B as an amateur model for Chinese-to-English (ZH-EN) on MQM23. $\bar{\textbf{Log(P)}}$ denotes mean log-probability across all target tokens.
Weighted Ensemble.   We investigate whether the averaged ensemble baseline used in our experiment is a sufficiently strong setup, or if a weighted combination of model probabilities could yield further improvements. Specifically, we evaluate a weighted ensemble approach that linearly combines the output probabilities of a larger and a smaller model, allowing us to examine whether tuning the weight parameter provides a meaningful advantage over uniform averaging.
@@
where $\gamma\in[0,1]$ is the weight factor that controls the contribution of each model. When $\gamma=0.5$, this formulation corresponds to that reported in our main results. To analyze the impact of $\gamma$, we conduct a parameter sweep on the summarization task using LLaMA (3B, 1B), varying $\gamma$ from 0 to 1 in increments of 0.05. The results, presented in Figure 3, show that the weighted ensemble achieves its highest performance at $\gamma=0.55$, which is nearly identical to the performance at $\gamma=0.5$ (i.e., the averaged ensemble baseline). This indicates that the simple averaged ensemble is already a strong and competitive configuration, and that further weighting does not lead to meaningful improvement.
Case Study.
To provide a qualitative analysis, Table 5 presents a case study comparing ContrastScore with a single expert model (Qwen 3B) and an amateur model (Qwen 0.5B) for Chinese-to-English (ZH-EN) machine translation. It can be observed that the expert model misranks Hypothesis 1 as the second-best translation, whereas human evaluators rank it as the best (Rank 1). This misranking is largely influenced by the expert model assigning a very low probability (0.000457) to “logo”, a key token in the translation. In contrast, the amateur model assigns a significantly higher probability (0.06592) to “logo”, demonstrating greater confidence in its relevance. ContrastScore leverages this probability discrepancy to adjust the probability of “logo”, thereby increasing the overall score of Hypothesis 1 and correctly ranking it as the best translation. This case illustrates how ContrastScore mitigates expert model underestimation of critical tokens, leading to improved evaluation alignment with human judgment.
In the case of Hypothesis 3, all methods (i.e., expert, amateur, and ContrastScore) correctly rank it as the worst translation (Rank 3). This consistency occurs because both the expert and amateur models assign low probabilities across all tokens, indicating poor translation quality. Unlike in Hypothesis 1, where significant probability discrepancies require adjustment, ContrastScore makes minimal modifications since the expert and amateur models are already in agreement in their evaluation. This demonstrates that ContrastScore does not introduce unnecessary changes when the models agree, ensuring that it only refines scores when meaningful corrections are needed.
@@ 7 Conclusion
In this paper, we introduce ContrastScore, an evaluation metric that leverages structured disagreement between two language models of differing capacities. Unlike conventional LLM-based evaluation methods that rely solely on a single model’s likelihood estimates, ContrastScore incorporates a weaker auxiliary model as a contrastive signal to adjust probability scores for better alignment with human judgments. Extensive experiments demonstrate that ContrastScore consistently achieves stronger correlation with human evaluations than both single-model and ensemble-based baselines. Furthermore, ContrastScore is computationally-efficient and can effectively mitigate likelihood bias, resulting in a more robust evaluation.
Limitations
This work presents the following potential limitations: (1) Our investigation is limited to the LLaMA and Qwen families of models. While ContrastScore demonstrates effectiveness within these two model families, further evaluation across a broader range of large language model architectures is necessary to establish its generalizability. (2) This work relies on token-level probabilities and therefore cannot combine models from different families due to differences in their subword tokenization. In future work, word-level probability could be explored to enable mixing across model families. (3) The metrics explored in this paper are not competitive to the state-of-the-art (SOTA) metrics developed for MT and summarization, such as MetricX (Juraska et al., 2024) or XCOMET-XXL (Guerreiro et al., 2024), which may use much larger models and/or fine-tuning on human annotations. We do not claim in this paper to beat SOTA metrics, but that contrastive principles, involving two models, are superior to non-contrastive principles for evaluation metric design within one evaluation paradigm (BARTScore, in our case). To design competitive metrics, future work should consider involving (considerably) larger models than we did in this work, beyond 8B parameters.
Acknowledgments
We thank the anonymous reviewers for their constructive feedback and helpful suggestions. Xiao Wang is supported by a studentship funded by the Department of Computer Science, University of Manchester.
Appendix A Datasets, Metrics and Prompts
A.1 Datasets
Machine Translation.
MQM22 and MQM23 datasets are annotated by professional translators from the WMT22 Freitag et al. (2022) and WMT23 Freitag et al. (2023) Metrics Shared Tasks, based on the Multidimensional Quality Metrics (MQM) framework Lommel et al. (2014). MQM22 covers 3 language pairs: English-German (EN-DE), Chinese-English (ZH-EN) and English-Russian (EN-RU), and comprises 1315, 1875, and 1315 segments for language pair EN-DE, ZH-EN, EN-RU respectively, with every segments including 15 system outputs. MQM23 covers 3 language pairs: English-German (EN-DE), Chinese-English (ZH-EN) and Hebrew-English (HE-EN), including 460, 1177, and 820 segments and 12, 15, and 12 systems for EN-DE, ZH-EN, HE-EN respectively.
Summarization.
SummEval Fabbri et al. (2021) contains 1600 model-generated summaries of CNN/DailyMail articles, each annotated by 3 experts and 5 crowd workers. It covers aspects of coherence, consistency, fluency, and relevance. QAGS-XSUM Wang et al. (2020) includes 239 system outputs from a fine-tuned BART on XSUM dataset, focusing on factuality aspect.
A.2 Baseline Metrics
BLEU Papineni et al. (2002) is based on the precision of n-grams between the MT output and its reference weighted by a brevity penalty.
CHRF Popović (2015) uses character n-grams instead of word n-grams to compare the MT output with the reference.
ROUGE  Lin (2004) measures the lexical overlap between the hypothesis and reference. We consider 3 variants ROUGE-1, ROUGE-2, and ROUGE-L.
COMET Rei et al. (2020) is a learnt metric that is fine-tuned to produce evaluation scores for a given translation by comparing its representation to source and reference embeddings.
COMET-KIWI Rei et al. (2022) is a reference-free variant of COMET for machine translation evaluation. It uses a multilingual encoder to take only the source and system output as input and predicts a continuous quality score.
BERTScore Zhang et al. (2020) leverages contextual embeddings from BERT to compare words in candidate and reference sentences using cosine similarity.
MoverScore Zhao et al. (2019) measures semantic similarity between a candidate and a reference by combining contextualized embeddings with Earth Mover’s Distance.
BARTScore Yuan et al. (2021) is a generative metric that uses BART to evaluate the generated text by calculating the probabilities of the tokens.
A.3 Prompts
The prompts for the summarization and machine translation tasks are presented in Table 6.
@@ A.3 Prompts
A.4 Additional Details
Artifact Licenses
We use the following artifacts:
  • Qwen 2.5 - Qwen License Agreement
  • LLaMA 3.1 and 3.2 - LLaMA 3.1 and 3.2 Community License Agreement
  • COMET22 - Apache 2.0 License
  • COMET-Kiwi - CC BY NC SA 4.0
  • SummEval - MIT License
  • WMT23 - MIT License
Our use of those artifacts complies with license terms and applicable intended use policies.
PII information in data
We did not specifically check whether the datasets used contain PII information. However, we have noticed that WMT23 authors made an effort to mask PII information with special tags. We do not disseminate any new dataset; therefore, PII protection is out of the scope of our work.
Package Versions and Hardware
We use ‘unbabel-comet‘ version 2.2.0 to run baselines for MT evaluation. Experiments were conducted on the university SLURM computing cluster with various available GPUs. For benchmarking purposes, we used H100 GPUs with 96GB of VRAM.
Appendix B Analysis
B.1 The impact of hyperparameter gamma
@@
Figure 2: Exploration of the impacts of $\gamma$. Testing correlation between evaluator score and human score in ZH-EN language pair on MQM23.
B.2 Weighted ensemble
To explore whether the advantages of ContrastScore can be replicated by a simpler method based on probability combination, we evaluate a weighted ensemble approach that linearly merges the output probabilities of the larger and smaller models. The results are presented in the Figure 3.
@@ Appendix C Results
C.1 Likelihood Bias
The detailed likelihood bias scores for every datasets of machine translation and summarization tasks are shown in Table 8 and Table 9. ContrastScore demonstrates consistent effectiveness in reducing likelihood bias across both machine translation and summarization tasks, particularly in challenging language pairs and critical evaluation aspects. In machine translation in Table 8, ContrastScore yields substantial improvements for the LLaMA family across diverse language pairs, especially for low-resource or morphologically complex pairs such as EN-RU and HE-EN. In summarization in Table 9, ContrastScore offers the most pronounced improvements in factuality and coherence, which are often the most difficult dimensions for language models. Overall, ContrastScore effectively mitigates likelihood bias compared to both single-model and ensemble baselines in both the LLaMA and Qwen families.
C.2 Efficiency
We use processing speed as an indicator of efficiency, with detailed results for both machine translation and summarization evaluation tasks presented in Table 7. Processing more samples per second indicates higher evaluation efficiency. These results demonstrate that ContrastScore offers substantial gains in evaluation efficiency, especially when leveraging smaller models.
LLaMA SUM MT Qwen SUM MT
Single
1B 63.29 339.04 0.5B 141.96 552.68
3B 47.61 153.25 3B 48.21 194.22
8B 26.43 72.17 7B 26.43 99.84
Ensemble
(3B,1B) 39.51 110.63 (3B,0.5B) 44.94 154.5
(8B,1B) 23.81 66.8 (7B,0.5B) 26.11 93.92
(8B,3B) 19.81 53.95 (7B,3B) 20.35 71.04
Contrast
(3B,1B) 39.12 109.31 (3B,0.5B) 44.06 153.99
(8B,1B) 23.33 66.11 (7B,0.5B) 25.71 93.41
(8B,3B) 19.05 53.45 (7B,3B) 19.61 70.97
Table 7: Processing Speed for Machine Translation and summarization evaluation tasks. Measured in Samples Per Second with batch size of 16 on single H100 GPU.
Settings MQM22 MQM23
EN-DE ZH-EN EN-RU AVG EN-DE ZH-EN HE-EN AVG
LLaMA Single 1B 0.437 0.312 0.276 0.342 0.142 0.237 0.257 0.212
3B 0.426 0.290 0.253 0.323 0.146 0.284 0.304 0.245
8B 0.367 0.287 0.238 0.297 0.181 0.309 0.565 0.352
Ensemble (3B,1B) 0.324 0.219 0.101 0.215 -0.002 0.169 0.202 0.123
(8B,1B) 0.278 0.189 0.074 0.180 0.006 0.148 0.302 0.152
(8B,3B) 0.298 0.227 0.140 0.222 0.079 0.223 0.385 0.229
Contrast (3B,1B) 0.195 0.143 -0.165 0.058 -0.106 0.096 0.088 0.026
(8B,1B) 0.164 0.131 0.016 0.104 -0.044 0.093 0.353 0.134
(8B,3B) 0.187 0.141 -0.011 0.106 -0.051 0.110 0.350 0.137
Qwen Single 0.5B 0.493 0.233 0.296 0.341 0.184 0.214 0.358 0.252
3B 0.580 0.367 0.379 0.442 0.279 0.476 0.598 0.451
7B 0.559 0.383 0.383 0.441 0.333 0.438 0.616 0.463
Ensemble (3B,0.5B) 0.463 0.275 0.256 0.331 0.115 0.370 0.498 0.327
(7B,0.5B) 0.421 0.236 0.202 0.286 0.095 0.304 0.483 0.294
(7B,3B) 0.490 0.303 0.280 0.358 0.206 0.376 0.562 0.381
Contrast (3B,0.5B) 0.432 0.257 0.233 0.307 0.087 0.350 0.505 0.314
(7B,0.5B) 0.415 0.239 0.206 0.287 0.092 0.296 0.518 0.302
(7B,3B) 0.395 0.235 0.187 0.272 0.091 0.292 0.499 0.287
Table 8: Likelihood bias of machine translation task for the LLaMA and Qwen family. ContrastScore can effectively mitigate the likelihood bias compared to both single and ensemble methods on MQM22 and MQM23 datasets.
Settings Q-XUM SummEval
Factuality Coherence Consistency Fluency Relevance AVG
LLaMA Single 1B 0.382 0.104 0.471 0.539 0.279 0.348
3B 0.289 0.169 0.477 0.573 0.321 0.385
8B 0.267 0.161 0.477 0.575 0.311 0.381
Ensemble (3B,1B) 0.249 0.095 0.406 0.488 0.240 0.308
(8B,1B) 0.225 0.119 0.417 0.507 0.261 0.326
(8B,3B) 0.242 0.152 0.450 0.538 0.295 0.359
Contrast (3B,1B) 0.233 -0.008 0.279 0.339 0.122 0.183
(8B,1B) 0.220 0.074 0.337 0.424 0.214 0.262
(8B,3B) 0.219 0.048 0.326 0.400 0.185 0.240
Qwen Single 0.5B 0.347 0.139 0.489 0.567 0.308 0.376
3B 0.349 0.173 0.492 0.570 0.332 0.392
7B 0.373 0.185 0.489 0.584 0.334 0.398
Ensemble (3B,0.5B) 0.282 0.170 0.461 0.567 0.318 0.379
(7B,0.5B) 0.289 0.194 0.464 0.582 0.332 0.393
(7B,3B) 0.345 0.151 0.472 0.560 0.295 0.369
Contrast (3B,0.5B) 0.294 0.017 0.364 0.410 0.152 0.236
(7B,0.5B) 0.318 0.077 0.411 0.488 0.208 0.296
(7B,3B) 0.353 0.112 0.430 0.526 0.249 0.329
Table 9: Likelihood bias of summarization task for the LLaMA and Qwen family. ContrastScore can effectively mitigate the likelihood bias compared to both single and ensemble methods on QAGS-XSUM and SummEval datasets.
Evaluators MQM22 MQM23 AVG
EN-DE ZH-EN EN-RU EN-DE ZH-EN HE-EN
LLaMA Single 1B 0.407 0.493 0.495 0.655 0.625 0.440 0.519
3B 0.415 0.497 0.498 0.661 0.629 0.441 0.524
8B 0.439 0.503 0.515 0.676 0.648 0.511 0.549
Ensemble (3B,1B) 0.416 0.498 0.503 0.665 0.633 0.442 0.526
(8B,1B) 0.426 0.504 0.518 0.674 0.646 0.489 0.543
(8B,3B) 0.430 0.504 0.515 0.675 0.647 0.490 0.544
Contrast (3B,1B) 0.435 0.502 0.509 0.674 0.639 0.438 0.533
(8B,1B) 0.449 0.507 0.528 0.682 0.649 0.496 0.552
(8B,3B) 0.445 0.509 0.530 0.683 0.652 0.496 0.553
Qwen Single 0.5B 0.395 0.500 0.505 0.650 0.637 0.451 0.523
3B 0.417 0.513 0.502 0.665 0.659 0.477 0.523
7B 0.427 0.516 0.517 0.665 0.655 0.494 0.539
Ensemble (3B,0.5B) 0.414 0.510 0.511 0.663 0.656 0.479 0.546
(7B,0.5B) 0.418 0.512 0.518 0.665 0.655 0.491 0.539
(7B,3B) 0.425 0.518 0.518 0.671 0.663 0.497 0.543
Contrast (3B,0.5B) 0.428 0.517 0.519 0.675 0.666 0.482 0.549
(7B,0.5B) 0.435 0.519 0.530 0.676 0.666 0.499 0.548
(7B,3B) 0.435 0.521 0.531 0.676 0.666 0.499 0.554
Table 10: Pairwise Accuracy of evaluators with human scores in machine translation. bold represents best overall scores, while underline represents best scores within each model group (Baseline, LLaMA, Qwen). Overall, ContrastScore outperforms single and ensemble methods, as well as baseline metrics for both LLaMA and Qwen families.